### 03 Apr 2026 v1

# Council Mode: Mitigating Hallucination and Bias in LLMs via Multi-Agent Consensus

Abstract
Large Language Models (LLMs), particularly those employing Mixture-of-Experts (MoE) architectures, have achieved remarkable capabilities across diverse natural language processing tasks. However, these models frequently suffer from hallucinations—generating plausible but factually incorrect content—and exhibit systematic biases that are amplified by uneven expert activation during inference. In this paper, we propose the Council Mode, a novel multi-agent consensus framework that addresses these limitations by dispatching queries to multiple heterogeneous frontier LLMs in parallel and synthesizing their outputs through a dedicated consensus model. The Council pipeline operates in three phases: (1) an intelligent triage classifier that routes queries based on complexity, (2) parallel expert generation across architecturally diverse models, and (3) a structured consensus synthesis that explicitly identifies agreement, disagreement, and unique findings before producing the final response. We implement and evaluate this architecture within an open-source AI workspace. Our comprehensive evaluation across multiple benchmarks demonstrates that the Council Mode achieves a 35.9% relative reduction in hallucination rates on the HaluEval benchmark and a 7.8-point improvement on TruthfulQA compared to the best-performing individual model, while maintaining significantly lower bias variance across domains. We provide the mathematical formulation of the consensus mechanism, detail the system architecture, and present extensive empirical results with ablation studies.
Keywords Large Language Models $\cdot$ Mixture-of-Experts $\cdot$ Hallucination Mitigation $\cdot$ Bias Reduction $\cdot$ Multi-Agent Consensus $\cdot$ Structured Synthesis
1 Introduction
The rapid advancement of Large Language Models (LLMs) has fundamentally transformed the landscape of natural language processing and artificial intelligence (Vaswani et al., 2017). Models such as GPT-5.4, Claude Opus 4.6, and Gemini 3.1 Pro have demonstrated unprecedented capabilities in reasoning, generation, and comprehension tasks. To scale model capacity without proportionally increasing computational costs during inference, the Mixture-of-Experts (MoE) architecture has emerged as a dominant paradigm (Shazeer et al., 2017; Fedus et al., 2022). By activating only a subset of expert modules for each input token, MoE models achieve remarkable efficiency while maintaining or exceeding the performance of dense counterparts.
Despite these advances, MoE-based LLMs are particularly susceptible to two critical failure modes: hallucination and systematic bias. Hallucination refers to the generation of text that is fluent and plausible but factually incorrect or unsupported by the input context (Ji et al., 2023). In MoE architectures, the sparse routing mechanism can exacerbate this problem through several pathways. First, the gating network may consistently route certain types of queries to experts that have internalized incorrect correlations during training, a phenomenon known as expert collapse (Zoph et al., 2022). Second, the limited activation of experts means that the model’s response to a given query is determined by only a fraction of its total parametric knowledge, increasing the likelihood of epistemic gaps. Third, the routing decisions themselves can introduce systematic biases, as the gating function may develop preferences that correlate with surface-level features rather than semantic content (Riquelme et al., 2021).
Traditional approaches to mitigating hallucinations include Retrieval-Augmented Generation (RAG), which grounds model outputs in external knowledge sources, and Reinforcement Learning from Human Feedback (RLHF), which aligns model behavior with human preferences. While effective to varying degrees, these methods fundamentally operate within the confines of a single model’s parametric memory and reasoning capabilities. Recent research has demonstrated that multi-agent debate and consensus mechanisms can significantly enhance reasoning quality and factual accuracy (Du et al., 2023; Liang et al., 2023; Chen et al., 2023). The key insight is that different models, trained on different data with different architectures, develop complementary knowledge representations and reasoning strategies. When these models are queried independently and their outputs are systematically compared, idiosyncratic errors tend to be identified and corrected through cross-verification.
In this paper, we introduce the Council Mode, a structured multi-agent consensus framework designed to mitigate hallucinations and biases in LLM outputs. Rather than relying on a single model—whether dense or MoE-based—the Council Mode queries $N$ architecturally diverse frontier models in parallel and employs a dedicated synthesis model to aggregate their responses through a formalized consensus protocol. The synthesis process explicitly categorizes the combined knowledge into four structured sections: consensus points, disagreements, unique findings, and a comprehensive analysis. This structured approach ensures that the final output is not merely a majority vote but a nuanced integration that preserves valuable minority insights while filtering out individual model hallucinations.
Our implementation is integrated into an open-source AI workspace (Wu, 2026) that serves as the experimental platform for this study. Through extensive experimentation, we evaluate the Council Mode against individual frontier models on established hallucination and reasoning benchmarks.
Our main contributions are as follows:
  1. 1.
    We propose the Council Mode, a three-phase multi-agent consensus framework (Triage $\rightarrow$ Parallel Expert Generation $\rightarrow$ Consensus Synthesis) that systematically reduces hallucination and bias in LLM outputs.
  2. 2.
    We formalize the mathematical framework for the consensus synthesis mechanism, including the triage function, parallel dispatch protocol, and structured aggregation.
  3. 3.
    We provide a complete open-source implementation and comprehensive empirical evaluation demonstrating significant improvements over single-model baselines.
  4. 4.
    We conduct ablation studies analyzing the contribution of each pipeline component and the effect of expert diversity on consensus quality.
2 Related Work
2.1 Hallucination in Large Language Models
Hallucination in LLMs has been extensively studied across multiple dimensions. Ji et al. (2023) provide a comprehensive taxonomy distinguishing between intrinsic hallucination (output contradicting the source) and extrinsic hallucination (output that cannot be verified from the source). The causes of hallucination are multifaceted, including training data noise, exposure bias during autoregressive generation, and the fundamental limitation of compressing world knowledge into finite parameters (Zhang et al., 2023).
In the context of MoE models, hallucination risks are amplified by the sparse activation pattern. Zoph et al. (2022) demonstrated that MoE models can develop routing biases where certain experts become over-specialized, leading to knowledge gaps when queries fall outside their narrow expertise. Furthermore, the inconsistency between expert routing during training and inference—particularly under policy updates in reinforcement learning—can induce abrupt shifts in the active parameter subspace, destabilizing output quality (DeepSeek-AI, 2025).
2.2 Multi-Agent Debate and Consensus
The use of multiple LLM agents to improve output quality has gained significant traction. Du et al. (2023) introduced multi-agent debate, where multiple instances of the same model argue and refine their answers through multiple rounds. Liang et al. (2023) extended this concept with divergent thinking mechanisms. Chen et al. (2023) formalized multi-agent consensus seeking via LLMs, drawing parallels to classical consensus protocols in distributed systems. More recently, Xu et al. (2025) demonstrated that multi-agent architectures can transform hallucination-prone outputs into consensus-based reliable results in software testing.
Our Council Mode differs from prior work in several key aspects: (1) we use heterogeneous models from different providers rather than multiple instances of the same model, maximizing cognitive diversity; (2) we employ a structured synthesis protocol that explicitly categorizes consensus, disagreement, and unique findings rather than simple majority voting; and (3) we integrate an intelligent triage mechanism to optimize computational resources by bypassing the full council for trivial queries.
2.3 Ensemble and Voting Methods
Classical ensemble methods in machine learning aggregate predictions from multiple models to reduce variance and improve robustness (Dietterich, 2000). In the LLM context, Wang et al. (2022) proposed self-consistency, which samples multiple reasoning paths from a single model and selects the most consistent answer. Nair-Kanneganti et al. (2025) extended this to voting ensembles across different LLMs. Our approach goes beyond simple voting by employing a synthesis model that performs deep semantic analysis of expert responses, preserving the reasoning chains and evidence that support each claim.
3 Methodology: The Council Architecture
The Council Mode operates through a three-phase pipeline, as illustrated in Figure 1. We formalize each phase mathematically and describe the implementation details.
Figure 1: The Council Pipeline Architecture. Phase 1 uses a lightweight triage classifier (Seed 2.0 Pro) to determine query complexity. Trivial queries are answered directly. Non-trivial queries proceed to Phase 2, where three architecturally diverse expert models (GPT-5.4, Claude Opus 4.6, Gemini 3.1 Pro) generate independent responses in parallel via Promise.all(). Phase 3 synthesizes the expert outputs into a structured four-section response through the consensus model.
3.1 Phase 1: Intelligent Triage
To optimize computational resources, a lightweight triage classifier first evaluates the complexity of the user query $q$. The triage function $T:\mathcal{Q}\rightarrow\{0,1\}$ determines whether a full council deliberation is necessary:
$T(q)=\begin{cases}0\text{ (Bypass)},&\text{if }C_{\text{client}}(q)=0\text{ and }C_{\text{server}}(q)=0\\ 1\text{ (Council)},&\text{otherwise}\end{cases}$ (1)
The triage employs a two-stage filter. The client-side classifier $C_{\text{client}}$ uses regex-based pattern matching to identify trivial queries such as greetings, acknowledgments, or ultra-short inputs:
$C_{\text{client}}(q)=\begin{cases}0,&\text{if }q\in\mathcal{P}_{\text{greet}}\text{ or }(|q|\leq\tau_{\text{len}}\text{ and }|q|_{\text{tok}}\leq\tau_{\text{tok}})\\ 1,&\text{otherwise}\end{cases}$ (2)
where $\mathcal{P}_{\text{greet}}$ is the set of greeting patterns, $\tau_{\text{len}}=18$ characters, and $\tau_{\text{tok}}=4$ tokens. The server-side classifier $C_{\text{server}}$ invokes a lightweight LLM call with minimal thinking budget ($\text{thinkingLevel}=\text{``minimal''}$, $\text{temperature}=0.3$) that outputs a binary decision with an optional direct answer.
3.2 Phase 2: Parallel Expert Generation
For non-trivial queries ($T(q)=1$), the system dispatches the prompt to $N$ distinct expert models simultaneously. Let $\mathcal{E}=\{E_{1},E_{2},\dots,E_{N}\}$ be the set of expert models. Each expert independently generates a response:
$R_{i}=E_{i}(q,\mathcal{H},\mathcal{W}_{i})\quad\forall i\in\{1,2,\dots,N\}$ (3)
where $\mathcal{H}$ is the conversation history and $\mathcal{W}_{i}$ represents the optional web search results retrieved by expert $i$. The parallel dispatch is implemented via asynchronous concurrent execution (Promise.all() in our JavaScript implementation), ensuring that the total latency is bounded by the slowest expert rather than the sum of all expert latencies:
$t_{\text{total}}=\max_{i\in\{1,\dots,N\}}t_{i}+t_{\text{synthesis}}$ (4)
In our implementation, $N=3$, utilizing models from three different providers to maximize architectural diversity, as detailed in Table 1.
Table 1: Expert model configuration in the Council pipeline. All models are accessed via their official APIs to ensure maximum stability and native streaming capabilities. Max Output refers to the token limit enforced in our experimental setup, not the model’s intrinsic maximum.
Expert Model ID Provider Context (K) Max Output
GPT-5.4 gpt-5.4 OpenAI 272 4,000
Claude Opus 4.6 claude-opus-4-6 Anthropic 1,000* 4,000
Gemini 3.1 Pro gemini-3.1-pro-preview Google 1,048 4,000
Seed 2.0 Pro (Synth.) doubao-seed-2-0-pro ByteDance 256 8,000
* Claude Opus 4.6 defaults to 200K but supports up to 1M tokens.
3.3 Phase 3: Consensus Synthesis
The synthesis model $S$ receives the query $q$ along with the set of expert responses $\{R_{1},R_{2},\dots,R_{N}\}$ and performs a structured aggregation. The synthesis function is defined as:
$O=S\big(q,\{R_{1},R_{2},\dots,R_{N}\},\mathcal{P}_{\text{synth}}\big)$ (5)
where $\mathcal{P}_{\text{synth}}$ is the synthesis prompt containing 16 mandatory rules that govern the output structure. The output $O$ is decomposed into four sections:
$O=\langle O_{\text{consensus}},O_{\text{disagree}},O_{\text{unique}},O_{\text{analysis}}\rangle$ (6)
Consensus Extraction.
The consensus section identifies factual claims $f$ that are supported by all $N$ experts:
$O_{\text{consensus}}=\{f\mid f\in\bigcap_{i=1}^{N}\mathcal{F}(R_{i})\}$ (7)
where $\mathcal{F}(R_{i})$ extracts the set of factual claims from response $R_{i}$.
Disagreement Identification.
The disagreement section captures claims where experts provide conflicting information:
$O_{\text{disagree}}=\{(f,f^{\prime})\mid f\in\mathcal{F}(R_{i}),f^{\prime}\in\mathcal{F}(R_{j}),f\perp f^{\prime},i\neq j\}$ (8)
where $f\perp f^{\prime}$ denotes that claims $f$ and $f^{\prime}$ are contradictory.
Unique Findings.
Claims that appear in only one expert’s response are preserved as unique findings:
$O_{\text{unique}}=\bigcup_{i=1}^{N}\left(\mathcal{F}(R_{i})\setminus\bigcup_{j\neq i}\mathcal{F}(R_{j})\right)$ (9)
Comprehensive Analysis.
The final section integrates all evidence into a coherent, fact-checked response that resolves disagreements and contextualizes unique findings.
3.4 Theoretical Analysis: Why Council Reduces Hallucination
We provide a probabilistic argument for why the Council Mode reduces hallucination. Let $p_{i}$ denote the probability that expert $E_{i}$ hallucinates on a given claim. Assuming independence between experts (justified by their architectural diversity), the probability that all $N$ experts hallucinate on the same claim is:
[…]
For $N=3$ experts with individual hallucination rates $p_{1}=0.18$, $p_{2}=0.16$, $p_{3}=0.19$ (empirically measured on HaluEval QA), the probability that all three hallucinate on the same claim is $0.18\times 0.16\times 0.19\approx 0.0055$, representing a 97.1% reduction compared to the worst individual model. In practice, the reduction is moderated by correlations between models (shared training data, similar architectures), but the consensus mechanism still provides substantial error correction.
4 Implementation
The Council Mode is implemented within an open-source AI workspace designed for cloud-native deployment on serverless infrastructure. The implementation is built with Next.js 16 (App Router), React 19, and communicates with the client via a custom Server-Sent Events (SSE) protocol that streams the real-time reasoning states of all parallel experts.
4.1 SSE Event Protocol
The Council engine broadcasts expert state transitions in real-time, enabling the client to display the progress of each expert independently. The protocol defines the following event types, as shown in Table 2.
Table 2: SSE event protocol for the Council engine. Events are streamed in real-time to provide transparency into the consensus-building process.
Event Trigger Payload
council_expert_states Stream init Array of expert configurations
council_expert_state State transition Expert key, phase, metadata
council_expert_result Expert completion Content, thinking, citations
council_summary_state Synthesis transition Phase indicator
text Streaming delta Synthesis text chunk
council_triage Triage bypass Direct answer content
citations Final citations Array of title-URL pairs
[DONE] Stream end
Expert phase transitions follow the sequence: pending $\rightarrow$ searching $\rightarrow$ thinking $\rightarrow$ done $|$ error. The synthesis phase transitions follow: pending $\rightarrow$ thinking $\rightarrow$ answering $\rightarrow$ done.
4.2 Context Window Utilization
Each expert operates independently with its own context window. The significant variation in context window sizes across models (272K–1,048K tokens for Council experts) is a feature rather than a limitation: it ensures that different experts can process different amounts of historical context, contributing to response diversity. Figure 2 illustrates the context window capacities of all models in the Council pipeline.
Figure 2: Context window capacity (in thousands of tokens) for all models evaluated in this study. The diversity in context sizes contributes to complementary knowledge retrieval across the Council pipeline and baseline models.
5 Experiments and Results
We evaluated the Council Mode against individual frontier models on multiple established benchmarks. All experiments were conducted through our open-source platform using official API integrations.
5.1 Experimental Setup
Benchmarks.
We evaluate on: (1) HaluEval (Li et al., 2023), a large-scale hallucination evaluation benchmark covering QA, summarization, and dialogue tasks; (2) TruthfulQA (Lin et al., 2021), which measures the tendency of models to generate truthful and informative answers; and (3) a custom Multi-Domain Reasoning benchmark consisting of 500 questions across six domains (Science, History, Medicine, Law, Technology, Finance) with varying complexity levels (1–10 reasoning steps).
Baselines.
We compare against five individual models: GPT-5.4, Claude Opus 4.6, Gemini 3.1 Pro, DeepSeek V3.2, and Seed 2.0 Pro, each evaluated independently under identical conditions. Note on Benchmark Data: To ensure a fair and consistent comparison, all benchmark scores reported in this paper (including those for individual baseline models) were obtained by running our own evaluation scripts in a unified environment via official APIs, rather than citing numbers from external leaderboards or official technical reports.
Metrics.
We report: (1) Hallucination Rate (%): the percentage of generated claims that are factually incorrect; (2) Truthful Score (%): the percentage of answers rated as truthful by the TruthfulQA evaluation framework; (3) Informative Score (%): the percentage of answers that are both truthful and informative; (4) Accuracy (%): correctness on the multi-domain reasoning benchmark; and (5) Bias Variance ($\sigma^{2}$): the variance of performance scores across domains, measuring consistency.
5.2 Hallucination Rate Reduction
Figure 3 presents the hallucination rates on the HaluEval benchmark across three task categories. The Council Mode achieves the lowest hallucination rate in all categories, with particularly pronounced improvements in the Summarization task (13.6% vs. 20.1% for the best individual model, Claude Opus 4.6).
Figure 3: Hallucination rates (%) on the HaluEval benchmark across QA, Summarization, and Dialogue tasks. Lower values indicate better performance. The Council Mode consistently achieves the lowest hallucination rates across all task categories.
Table 3 provides detailed numerical results. The Council Mode achieves an average hallucination rate of 10.7%, compared to 16.7% for the best individual model (Claude Opus 4.6), representing a 35.9% relative reduction.
Table 3: Hallucination rates (%) on the HaluEval benchmark. Bold indicates best performance. $\Delta$ shows the relative reduction compared to the best individual model.
Model QA Summ. Dialogue Avg. $\Delta$
GPT-5.4 18.7 23.4 15.3 19.1
Claude Opus 4.6 16.2 20.1 13.8 16.7
Gemini 3.1 Pro 19.4 24.8 16.1 20.1
DeepSeek V3.2 21.3 26.5 18.7 22.2
Seed 2.0 Pro 22.8 28.2 19.5 23.5
Council Mode 10.1 13.6 8.4 10.7 -35.9%
5.3 TruthfulQA Results
Figure 4 shows the TruthfulQA benchmark results. The Council Mode achieves a Truthful score of 82.6% and an Informative score of 91.3%, surpassing the best individual model (Claude Opus 4.6 at 74.8% Truthful) by 7.8 percentage points.
Figure 4: TruthfulQA benchmark results showing Truthful (%) and Informative (%) scores. The Council Mode (highlighted) achieves the highest scores in both metrics.
5.4 Bias Mitigation Analysis
To assess bias mitigation, we measured the distribution of factual consistency and neutrality scores across 500 test prompts spanning six domains. Figure 5 presents the scatter plot of individual model outputs versus the Council Mode output.
Figure 5: Scatter plot of Factual Consistency Score vs. Neutrality Score across 500 test prompts. Individual models (circles, triangles, squares) show wide dispersion, while the Council Mode (diamonds) exhibits a tight cluster with high consistency and neutrality, indicating effective bias mitigation.
The key observation is that individual models exhibit wide variance in both dimensions, reflecting domain-specific biases. The Council Mode outputs form a tight cluster in the high-consistency, high-neutrality region, with a bias variance of $\sigma^{2}=0.003$ compared to $\sigma^{2}=0.021$$0.028$ for individual models—a reduction of approximately 85–89%.
5.5 Domain-Specific Hallucination Analysis
Figure 6 presents a heatmap of hallucination rates across six knowledge domains. The Council Mode achieves the lowest hallucination rate in every domain, with particularly strong performance in Law (6.5%) and History (7.1%), where individual models struggle most with factual accuracy.
Figure 6: Heatmap of hallucination rates (%) across six knowledge domains. The Council Mode (bottom row) consistently achieves the lowest rates, with the most pronounced improvements in domains where individual models exhibit the highest error rates.
5.6 Performance Scaling with Task Complexity
Figure 7 illustrates how accuracy scales with increasing task complexity, measured by the number of reasoning steps required. While all models show declining accuracy as complexity increases, the Council Mode demonstrates significantly more graceful degradation. At the highest complexity level (10 reasoning steps), the Council Mode maintains 71.2% accuracy compared to 50.8% for GPT-5.4 and 43.5% for DeepSeek V3.2, representing a 20.4–27.7 percentage point advantage.
Figure 7: Accuracy (%) as a function of task complexity (number of reasoning steps). The Council Mode (solid red line) maintains significantly higher accuracy than individual models (dashed lines) as complexity increases, with the performance gap widening at higher complexity levels.
5.7 Latency–Quality Trade-off
A natural concern with multi-model architectures is increased latency. Figure 8 presents the latency–quality trade-off. While the Council Mode incurs higher average latency (8.4s vs. 2.8–5.6s for individual models) due to the synthesis phase, the quality improvement (91.7% vs. 71.4–81.5%) represents a favorable trade-off for applications where accuracy is prioritized over speed.
Figure 8: Latency (seconds) vs. Quality Score (%) trade-off. The Council Mode achieves significantly higher quality at the cost of moderately increased latency. The parallel dispatch ensures that latency scales with the slowest expert rather than the sum.
5.8 Ablation Study
To understand the contribution of each component, we conducted ablation experiments, summarized in Table 4.
Table 4: Ablation study on the Council pipeline components. Each row removes one component from the full pipeline. Results are averaged across all HaluEval tasks.
Configuration Halluc. Rate Truthful Quality Latency
Full Council (3 experts + synthesis) 10.7% 82.6% 91.7% 8.4s
w/o Triage (always full council) 10.7% 82.6% 91.7% 12.1s
w/o Structured Synthesis (majority vote) 14.2% 77.3% 85.4% 6.8s
2 Experts + Synthesis 12.8% 79.1% 88.2% 7.1s
Same-model Ensemble ($3\times$ GPT-5.4) 15.6% 75.8% 83.1% 7.9s
Best Single Model (Claude Opus 4.6) 16.7% 74.8% 81.5% 4.1s
Key findings from the ablation study:
  1. 1.
    Triage primarily affects latency (reducing average latency by 30.6%) without impacting quality, confirming its role as a computational optimization.
  2. 2.
    Structured Synthesis is critical: replacing it with simple majority voting increases the hallucination rate by 32.7%, demonstrating that the synthesis model’s ability to reason about expert disagreements is essential.
  3. 3.
    Expert Diversity matters: using three instances of the same model (same-model ensemble) yields only modest improvements over the single model, confirming that architectural diversity is key to the Council’s effectiveness.
  4. 4.
    Number of Experts: reducing from 3 to 2 experts increases the hallucination rate by 19.6%, suggesting that three experts provide a good balance between diversity and computational cost.
6 Discussion
6.1 Why Heterogeneous Experts Outperform Homogeneous Ensembles
The ablation study reveals that the same-model ensemble ($3\times$ GPT-5.4) achieves an 18.3% relative reduction in hallucination compared to the single GPT-5.4 model, while the heterogeneous Council achieves 35.9%. This disparity arises because models from the same provider share training data, architectural biases, and failure modes. When GPT-5.4 hallucinates on a claim, another instance of GPT-5.4 is likely to produce the same hallucination. In contrast, Claude Opus 4.6 and Gemini 3.1 Pro, trained on different corpora with different architectures, are more likely to provide correct information or at least different errors, enabling the synthesis model to identify and correct the hallucination.
6.2 The Role of Structured Synthesis
Simple majority voting treats all claims as binary (correct/incorrect) and discards valuable reasoning chains. The structured synthesis approach preserves the full context of each expert’s response, enabling the synthesis model to: (1) identify why experts disagree, not just that they disagree; (2) assess the strength of evidence supporting each claim; and (3) preserve unique insights that might be lost in a voting scheme. This is particularly important for complex, multi-faceted questions where the “correct” answer is nuanced.
6.3 Limitations
[…]
7 Conclusion
In this paper, we presented the Council Mode, a multi-agent consensus architecture designed to mitigate hallucinations and biases inherent in traditional LLMs and MoE models. By leveraging the cognitive diversity of multiple heterogeneous frontier models and employing a structured synthesis pipeline that explicitly identifies consensus, disagreement, and unique findings, the Council Mode delivers highly reliable and factually grounded responses. Our comprehensive evaluation demonstrates significant improvements across multiple benchmarks: a 35.9% relative reduction in hallucination rates on HaluEval, a 7.8-point improvement on TruthfulQA, and an 85–89% reduction in bias variance across domains.
The open-source implementation provides a robust foundation for future research into multi-model orchestration. We believe that the Council paradigm represents a promising direction for building trustworthy AI systems, complementing rather than replacing ongoing efforts in model-level alignment and safety.
Acknowledgements
We extend our deepest gratitude to our academic advisor, Dr. Yanna Feng, for her invaluable guidance and intellectual support throughout this research.
We are especially grateful to our testing and support team for their rigorous evaluation and constructive feedback, which were instrumental in refining the Council architecture: Bolun Liu (M.S.), Weilin Cai (M.S.), Xinwei Du (M.S.), Zihao Su (B.S.), and Ran Wang (B.S.).
Furthermore, we would like to express our profound appreciation to the parents of Shuai Wu, Rongbao Wu and Xiaohua Gao, for their unwavering personal support and encouragement, which made this endeavor possible.
This work was supported by the open-source community. The source code is available at https://github.com/Noah-Wu66/Vectaix-AI under the MIT License.
Appendix A Synthesis Prompt Template
The synthesis model receives a system prompt containing 16 mandatory rules. The key rules include:
  1. 1.
    The output must contain exactly four sections: Model Consensus, Model Disagreement, Unique Findings, and Comprehensive Analysis.
  2. 2.
    Each consensus point must be supported by evidence from all three experts.
  3. 3.
    Disagreements must include the reasoning behind each expert’s position.
  4. 4.
    The Comprehensive Analysis must directly answer the user’s question, integrating all evidence.
  5. 5.
    Citations from expert responses must be preserved and attributed.
  6. 6.
    The synthesis must not introduce new claims not present in any expert response.
  7. 7.
    When experts unanimously agree, the synthesis should present the claim with high confidence.
  8. 8.
    When experts disagree, the synthesis should present all perspectives and indicate the weight of evidence.
Appendix B Complete Benchmark Results
Table 5 presents the complete benchmark results across all evaluation dimensions.
Table 5: Complete benchmark results across all evaluation dimensions. Best results in bold.
Model HaluEval TruthfulQA Reasoning Bias $\sigma^{2}$ Latency Quality
Avg. $\downarrow$ Truthful $\uparrow$ Acc. $\uparrow$ $\downarrow$ (s) Score $\uparrow$
GPT-5.4 19.1 71.2 72.4 0.023 3.2 78.3
Claude Opus 4.6 16.7 74.8 74.1 0.021 4.1 81.5
Gemini 3.1 Pro 20.1 68.5 70.8 0.028 2.8 76.2
DeepSeek V3.2 22.2 65.3 67.5 0.025 5.6 73.8
Seed 2.0 Pro 23.5 63.1 65.2 0.027 3.8 71.4
Council Mode 10.7 82.6 83.9 0.003 8.4 91.7